AB INITIO STRUCTURES, VIBRATIONS, AND THERMOCHEMISTRY OF IO, IO^{−}, AND IO^{+}

\underline{P. HASSANZADEH} ^{a} AND K. K. IRIKURA, Physical and Chemical Properties Division, National Institute of Standard and Technology, Gaithersburg, MD 20899.

Ab initio internuclear distances and anharmonic constants for IO, IO^{−}, and IO^{+} were calculated at the CCSD(T) level using 6-311+G(3df,2p) basis sets. The reliability of the calculations was tested against the available experimental values for IO and IO^{−}. Electronic energies were also calculated at the CCSD(T) level using the 6-311+G(3df,2p) basis sets, Sadlej’s polarized basis sets (PBS), and Wadt and Hay’s effective core potential (ECP) basis sets at the experimental geometries and corrected by the experimental zero-point vibrational energies. The enthalpy of formation of the hypoiodite anion IO^{−} (^{1}Σ^{+}) (Δ_{f}H_{0}^{o} -109 ± 5 kJ/mol, uncertainties are 2σ) was obtained from ab initio energetics for the metathesis reactions IO^{−} + Cl^{−} → ClO^{−} + I^{−} and IO^{−} + Cl_{2} → ClO^{−} + ICl. The reliability of the computational procedure was tested by performing analogous calculations for BrO^{−} and comparing the results with the accepted value for the enthalpy of formation of BrO^{−}. The enthalpy of formation for IO (X^{2}Π_{3/2} (Δ_{f}H_{0}^{o} = 120 ± 5 kJ/mol) was deduced from the calculated Δ_{f}H_{0}^{o}(IO^{−}) and the experimental electron affinity of IO (EA = 229.43 ± 0.06 kJ/mol) measured by Gilles et al.. The enthalpy of formation of IO^{+}(^{3}Σ^{−}) (Δ_{f}H_{0}^{o} = - 1060 ± 5 kJ/mol) was also derived from the present enthalpy of formation of IO and the experimental adiabatic ionization energy of IO (IE = 939.4 ± 1.7 kJ/mol) measured by Zhang et al. The corresponding bond dissociation energies are D_{0}^{o}(IO)= 167 ± 5 kJ/mol, D_{0}^{o}(IO) = 234 ± 5 kJ/mol, and D_{0}^{o}(IO^{+})= 303 ± 5 kJ/mol. JANAF-style thermodynamic functions for IO, IO^{−}, and IO^{+} from 0 to 1000 K are also tabulated.

aGuest Researcher